### v3

# Ultralight Black Holes as Sources of High-Energy Particles

Abstract
The memory burden effect, stating that the amount of information stored within a system contributes to its stabilization, is particularly significant in systems with a high capacity for information storage such as black holes (BHs). In these systems, the evaporation process is halted, at the latest, after approximately half of the BH’s initial mass has been radiated away. Consequently, light primordial BHs (PBHs) of mass $m_{\rm PBH}\lesssim 10^{15}$ g, which are expected to have fully evaporated by present time, may remain viable candidates for dark matter (DM). In this scenario, we demonstrate that their mergers would continue to occur today, leading to the formation of “young” BHs that resume evaporating, producing ultrahigh-energy cosmic rays detectable by current experiments. The emission spectrum would be thermal in all Standard Model particle species, offering a clear and distinguishable signature. Current measurements of the isotropic neutrino flux at Earth are in tension with light PBHs as DM candidates within the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$. Contrarily to existing bounds, our results depend uniquely on the mass of the remnant and not on the specific model building details of the stabilized phase. We also discuss the potential for refining these constraints through gamma-ray and cosmic-ray observations, as well as gravitational wave detections.
Introduction — Black holes (BHs), predicted by general relativity, are not only fascinating astrophysical objects but also intriguing sources of particle emission. A BH of mass $m_{\rm BH}$ emits high-energy particles lighter than its Hawking temperature, $T_{H}=(8\pi Gm_{\rm BH})^{-1}$ Hawking (1975). In the semiclassical regime where quantum gravity effects are neglected throughout their evaporation, the mass loss rate for an emitting BH follows $\dot{m}_{\rm BH}\propto m_{\rm BH}^{-2}$. The Hawking radiation has been searched for in terms of the BH emission into photons Carr et al. (2016); Boudaud and Cirelli (2019); DeRocco and Graham (2019); Laha (2019); Laha et al. (2020); Chan and Lee (2020), neutrinos Dasgupta et al. (2020); Bernal et al. (2022a), gravitons Hooper et al. (2020), and exotic particles Hooper et al. (2019); Schiavone et al. (2021); Auffinger (2023); Baker and Thamm (2023).
However, it has been suggested that the information contained within a BH affects its decay, slowing down the evaporation process through a “memory burden” effect Dvali (2018a), which becomes significant after a portion of its mass has evaporated Dvali et al. (2020). The timescale that signals the breakdown of the semiclassical description, which is also the latest the memory burden effect sets in, reads
$\tau_{\rm sc}\simeq(1-\epsilon^{3})\,\tau_{\rm evap}\,,$ (1)
where $\tau_{\rm evap}$ is the BH evaporation time in the semiclassical limit and the prefactor accounts for a fraction $1-\epsilon$ of the BH mass being lost in the process. Here, we fix $\epsilon=1/2$. The subsequent dynamics of the memory-burdened BH for times $t\gtrsim\tau_{\rm sc}$ remain uncertain due to the influence of quantum gravity effects Dvali et al. (2020). One possible outcome is that the configuration becomes unstable and eventually breaks down. Alternatively, evaporation could persist at a reduced rate, controlled by powers of the gravitational coupling, $\alpha_{\rm gr}\simeq S^{-1}$. In what follows, we assume the former scenario as the likely outcome indicated by theoretical analyses Dvali et al. (2020, 2024a).
When quantum gravity effects are taken into account, the BH evaporation rate is slowed down and the BH lifetime is extended as Dvali et al. (2020)
$\tau\simeq\tau_{\rm sc}\,S^{1+k}\,,$ (2)
where $S=4\pi Gm_{\rm BH}^{2}$ represents the BH entropy. Theoretical models such as the holographic description of BHs Dvali (2018b) and the microscopic description in terms of marginally bound gravitons—the so-called BH $N$-portrait Dvali and Gomez (2013a, 2014a, 2014b)- independently suggest that $k=1$ Dvali et al. (2024a), while numerical analyses give $k\geq 1$ Dvali et al. (2024a). In what follows, we will derive bounds independent of this quantity.
The effects described above are particularly relevant for BHs that have formed primordially, or PBHs. A PBH emitting semiclassically has not fully evaporated to date if its mass is $m_{\rm PBH}\gtrsim 10^{15}$ g. As a result, PBHs have long been considered potential dark matter (DM) candidates Zel’dovich and Novikov (1967); Hawking (1971); Carr and Hawking (1974) (see e.g., Escrivà et al. (2022) for a review). Even if PBHs constitute only a fraction $f_{\rm PBH}$ of the present DM abundance, they could still reveal their existence through a variety of astrophysical Pani and Loeb (2014); Boucenna et al. (2018); Smyth et al. (2020); Carr et al. (2021a, b); Saha and Laha (2022); Tran et al. (2024); Saha et al. (2024); Thoss and Burkert (2024), particle Lunardini and Perez-Gonzalez (2020), and cosmological phenomena Kohri and Yokoyama (2000); Baumann et al. (2007); Carr et al. (2010); Hook (2014); Fujita et al. (2014); Hooper and Krnjaic (2021); Perez-Gonzalez and Turner (2021); Smyth et al. (2022); Mazde and Visinelli (2023); Bernal et al. (2022b); Calabrese et al. (2023, 2024); Gehrman et al. (2023); Borah et al. (2024); Boccia et al. (2024).
The memory burden effect introduces a new mass window for PBHs as DM candidates below $10^{15}$ g, with BHs of mass around $10^{8}$ g shown to be viable while satisfying all known constraints Dvali et al. (2020). For $k=1$, Eq. (2) predicts that memory-burdened BHs as light as $3\times 10^{3}\rm g$ can be DM candidates. A viable mechanism for their production within this mass range $3\times 10^{4}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$, here dubbed the ultralight BH window, was introduced in Ref. Dvali et al. (2021). A more precise mass range for memory-burdened PBHs constituting DM has been recently derived Alexandre et al. (2024); Thoss et al. (2024). Other consequences of memory-burdened BHs have also been recently explored Balaji et al. (2024); Haque et al. (2024); Riotto and Silk (2024); Barman et al. (2024a, b); Bhaumik et al. (2024); Kohri et al. (2024); Jiang et al. (2024); Chianese et al. (2024); Athron et al. (2024).
In this Letter, we discuss the observational consequences of memory-burdened PBHs as DM. We argue that these PBHs, locked in binary configurations, merge to form new BHs of similar mass. Upon merging, the new object is a young BH, no longer subject to the memory burden effect. As such, the freshly-formed BH starts emitting again for a duration $\tau_{\rm sc}$, at much lower redshifts than their progenitors, thus leading to novel signatures that are unique to the memory burden phenomenon.
Memory-burdened PBH mergers could potentially occur within our own Galaxy if these objects constitute a significant DM fraction. The light BHs resulting from their mergers would be bright sources of gamma and cosmic rays, detectable at observatories Coleman et al. (2023), as well as sources of high and ultra high-energy neutrinos detectable by neutrino telescopes Abbasi et al. (2010, 2011). Gravitons in ultra-high frequency windows are also emitted, providing new target opportunities for detection facilities. Here, we work with natural units $\hbar=c=1$.
Memory burden effect — Saturated configurations offer valuable insights into objects with high information-storage capacity, revealing connections between gravitational and non-gravitational systems. Notably, “saturons”, configurations of maximal entropy constrained by the unitarity bound Dvali (2021a), exemplify these objects. BHs are a prime example of saturated configurations, as they obey the entropy-area law Bekenstein (1973). Saturated configurations also arise in non-gravitational theories, such as renormalizable field theories Dvali (2021a, b, c). Saturated configurations not only follow an area law, but also share several properties with BHs, such as thermal-like evaporation Dvali and Sakhelashvili (2022); Dvali et al. (2022a), the presence of an information horizon in the semiclassical regime Dvali (2021b, c, a, d); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), a timescale for information retrieval corresponding to the half-time of evaporation Dvali (2021a); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), and a maximal spin bound that cannot exceed their entropy, similar to the extremality condition for spinning BHs Dvali et al. (2022b, 2024b). Many features traditionally associated with BHs may arise from the broader behavior of saturated configurations, rather than being unique to gravity.
This framework provides new approaches for studying BHs from both theoretical and phenomenological perspectives. By analyzing the non-gravitational counterparts of BHs, we can use the calculability of renormalizable theories to better understand their enigmatic properties. Extending the study of saturated configurations to BHs allows us to predict new phenomenological signatures. For example, rotating saturons provide a microscopic description of maximally spinning BHs as graviton condensates with vorticity Dvali et al. (2022b), offering a topological explanation for the absence of Hawking radiation. This could lead to observable deviations in the gravitational radiation from BH mergers Dvali et al. (2024b).
Furthermore, the memory burden effect, studied in calculable saturated configurations, is expected to manifest in BHs as well Dvali et al. (2020, 2024a). The memory stored in a BH can be represented as $M$ qubits, corresponding to the occupation number in the flavor space of Goldstone bosons arising from broken symmetries Dvali (2021a, b, c). The number of degenerate states is $n_{\rm st}=2^{M}$, leading to an entropy $S=\log n_{\rm st}\simeq M$. As BHs decay, the increasing mass gap of the Goldstone bosons halts evaporation by the half-mass decay time. [^1] The memory burden effect is an explicit realization of the so-called quantum-breaking of classical configurations Dvali et al. (2013), discussed within the framework of the BH $N$-portrait Dvali and Gomez (2013b, 2014b), which questions the validity of the classical description of BHs and finds that it is violated around $\tau_{\rm sc}$. This can be modeled using the probability distribution for memory-occupied modes Dvali et al. (2024a),
[…]
where $N_{\rm G}$ represents the order parameter setting the timescale of quantum backreaction Dvali et al. (2024a). The probability distribution in (3) leads to a natural spread in the mass distribution of PBHs Dvali et al. (2024a), offering a unique signature and further insights into BH evolution and mergers. A careful dynamical analysis reveals that an evaporating BH asymptotically evolves into a memory-burdened BH with around half of its initial mass Dvali et al. (2024a). While it is possible for a BH with empty memory to exist, allowing for its complete evaporation without memory burden, such cases are exponentially rare, as indicated by the probability distribution in Eq. (3).
As suggested by the study of mergers Dvali et al. (2024b) of memory-burdened configurations Dvali et al. (2024a), only the merging point is affected by the memory burden, potentially emitting additional radiation, while the merged object is unaffected. Mapped to the gravitational counterpart, these results imply that the memory burden of the two merging BHs leads to significant deviations in the gravitational wave signal only at ringdown, when a new classical BH forms with a memory pattern governed by Eq. (3). Besides, the newly formed BH has an area, and thus an entropy, approximately four times larger than the combined areas of the two initial BHs, providing considerable memory to store all the non-emitted information. Therefore, post-merger the BH emits semiclassically for a time $\tau_{\rm sc}$ before stabilizing again. [^2] The merger of extremal ultralight PBHs was discussed in Ref. Bai and Orlofsky (2020) where it was argued that Hawking emission is reactivated. Indeed, the memory burden effect shares similarities with BHs that are semiclassically stabilized by extremality Dvali et al. (2024a). This leads to important phenomenological consequences, which are discussed next.
The cosmic ray flux from evaporating PBHs — The present fraction of memory-burdened PBHs formed at temperature $T_{f}$ is related to their primordial abundance $\beta$ through the conservation of the PBH comoving number, as
$f_{\rm PBH}\approx 10^{21}\,\beta^{\prime}\,\left(10^{8}{\rm\,g}/m_{\rm PBH}\right)^{1/2}\,.$ (4)
where $\beta^{\prime}=\gamma^{1/2}\,(0.7/h)^{2}\,(106.75/g_{*,f})^{1/4}\beta$ is the rescaled fraction in terms of the collapsed PBH fraction $\gamma$ Carr (1975), the reduced Hubble rate $h$, and the degrees of freedom at formation $g_{*,f}$. This expression generally does not apply to PBHs with masses $m_{\rm PBH}\lesssim 10^{15}$ g, as they have already evaporated and no longer contribute to today’s DM fraction $f_{\rm PBH}$. For these lighter PBHs, constraints are typically given in terms of the abundance at formation $\beta$. Here, we recast these bounds in terms of $f_{\rm PBH}$, reflecting the persistence of PBHs to present day.
The differential rate of PBH binary mergers describes a distribution of PBHs that decoupled from the Hubble flow before matter-radiation equality Ali-Haïmoud et al. (2017); Raidal et al. (2019); Liu et al. (2019). For a monochromatic mass function, the most conservative merger rate is Vaskonen and Veermäe (2020); Hütsi et al. (2021); Jedamzik (2020); Young and Hamers (2020); Jedamzik (2021)
$R_{\rm PBH}(t)=\frac{4.9\times 10^{16}}{\rm kpc^{3}\,yr}\,f_{\rm PBH}^{\frac{53}{37}}\,\left(\frac{t_{0}}{t}\right)^{\frac{34}{37}}\,\left(\frac{m_{\rm PBH}}{\rm g}\right)^{-\frac{32}{37}}S\,,$ (5)
where the suppression factor $S$ is the product of two components: a term $S_{1}\approx 0.24$ which accounts for the interactions between the binary system and the surrounding DM inhomogeneities, as well as neighboring BHs Hütsi et al. (2021), and $S_{2}(x)\approx\min\left[1,9.6\times 10^{-3}x^{-0.65}\exp\left(0.03\ln^{2}x\right)\right]$ parameterising the suppression caused by BHs absorbed by collapsed PBH clusters. Here, $x\equiv\left(t/t_{0}\right)^{0.44}f_{\mathrm{PBH}}$. The function $S_{2}(x)$ attains a minimum at $10^{-2}$, consistently with numerical simulations Inman and Ali-Haïmoud (2019); Franciolini et al. (2022). For light PBHs, dynamical captures induced by gravitational waves (GWs) as well as late-time dynamical capture are negligible Franciolini et al. (2022). However, local non-Gaussianity in primordial curvature perturbations can cluster PBHs at formation, potentially enhancing the merger rate by up to $\mathcal{O}(10^{7})$ Raidal et al. (2017); De Luca et al. (2021); Franciolini et al. (2022). Although some level of clustering is expected from various mechanisms at the formation of the PBH binary, we consider the most conservative rate expressed by Eq. (5) as the fiducial model.
[…]
The emission of newly formed BHs in Eq. (1) occurs on shorter timescales than astrophysical processes. For this, we estimate the galactic flux of particles from recently produced BHs from PBH mergers as
$\frac{{\rm d}\Phi_{i}}{{\rm d}E}\bigg{|}_{\rm gal}\simeq\tau_{\rm sc}\,\int\frac{{\rm d}\Omega}{4\pi}\int{\rm d}s\,R_{\rm PBH}\,\delta(r(s,\theta))\,\frac{{\rm d^{2}}N_{i}(E)}{{\rm d}E\,{\rm d}t}\,,$ (6)
where $\delta(r)\equiv\rho_{\rm DM,gal}(r)/\rho_{\rm DM}$ describes the Galactic enhancement in the DM density distribution Pujolas et al. (2021), which is here modeled as a Navarro-Frenk-White (NFW) profile Navarro et al. (1997), $\rho_{\rm DM,gal}(r)=\rho_{s}(r_{s}/r)(1+r/r_{s})^{-2}$, where we fix $r_{s}=15.6$ kpc and $\rho_{s}$ is taken so that the DM density in the vicinity of the Solar System is $\rho_{\odot}=0.45{\rm\,GeV\,cm^{-3}}$, while $r_{\rm vir}$ is defined as $\delta(r_{\rm vir})=200$. The function $r(s,\theta)$ accounts for the line-of-sight distance $s$ of the merger from Earth, with the integration over $s$ extending up to the value for which $r=r_{\rm vir}$. Averaging the differential flux over the angular direction $\theta$ and integrating in the energy bins yields the integral flux $\Phi_{i}^{\rm gal}(E)$. The extra-galactic (eg) contribution to the differential flux of the species $i$ is
$\frac{{\rm d}\Phi_{i}}{{\rm d}E}\bigg{|}_{\rm eg}\simeq\frac{\tau_{\rm sc}}{4\pi}\int_{0}^{z_{f}}\!{\rm d}z\left|\frac{{\rm d}t}{{\rm d}z}\right|R_{\rm PBH}(t(z))\frac{{\rm d^{2}}N_{i}(E^{\prime})}{{\rm d}E\,{\rm d}t}e^{-\eta_{i}(z)}\,,$ (7)
where ${\rm d}t/{\rm d}z$ is the cosmological line element, $E^{\prime}=E\,(1+z)$ accounts for the redshift of the energy $E^{\prime}$ at emission and $z_{f}$ is the redshift of PBHs at formation. The opacity $\eta_{i}$ describes the attenuation between $z$ and today at energy $E$.
Results — Figure 1 shows the integrated gamma-ray flux $\Phi_{i}^{\rm gal}(E)$ in Eq. (6), produced by newly-formed BHs from PBH mergers and for $f_{\rm PBH}=1$, for different values of the PBH masses and the conservative merger rate in Eq. (5). We have not included the eg contribution since, at energies above the TeV scale, the Universe is opaque to gamma rays due to gamma-gamma absorption (see Cao et al. (2023) for a review). Results are compared with the upper limits at 95% confidence level (CL) from the Pierre Auger Observatory (Auger) Savina (2021); Abreu et al. (2022, 2023), the Karlsruhe Shower Core and Array Detector (KASCADE), and KASCADE-Grande Apel et al. (2017). Lighter BHs are brighter due to the higher Hawking temperature and smaller entropy. For this, only the heavier BHs evade the bounds from the measured flux of gamma rays. We have not included any effects from physics beyond the SM, which would impact our results by altering the semiclassical evaporation timescale and the emission from the secondary spectrum.
Figure 1: The integrated gamma-ray flux emitted by newly-formed BHs from memory-burdened Galactic PBH mergers, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$. Also shown are the upper limits at 95% CL from Auger Savina (2021); Abreu et al. (2022, 2023), KASCADE, and KASCADE-Grande Apel et al. (2017).
We now discuss the emission of energetic neutrinos of all flavors by the same sources. The differential neutrino flux per solid angle and energy is reported by the IceCube collaboration using the cascades collected over 7.5 years of operation at energies above 60 TeV Abbasi et al. (2021a); Aartsen et al. (2020). Above the PeV, an upper limit at 95% CL on the neutrino flux is reported Aartsen et al. (2018, 2019). In Fig. 2 we compare these experimental results with the the differential flux of all-flavor neutrinos sourced by newly-formed BHs. We include the Galactic (continuous) and extragalactic (dashed) contributions assuming the merger rate in Eq. (5).
Figure 2: The differential flux of all-flavor neutrinos sourced by newly-formed BHs, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$. We show the Galactic (continuous) and the extra-galactic (dashed) contributions. The theoretical flux is compared with the all-flavor neutrino flux above 60 TeV Abbasi et al. (2021a); Aartsen et al. (2020) and with the upper bound for ultra-high energy neutrinos Aartsen et al. (2018, 2019).
Discussion — Assuming that a fraction $f_{\rm PBH}$ of the DM relic abundance consists of light PBHs, we derive a constraint on the model in the $(m_{\rm PBH},f_{\rm PBH})$ plane by performing a least-square analysis that accounts for the neutrino flux measurements and upper bounds reported by the IceCube collaboration Abbasi et al. (2021a); Aartsen et al. (2020, 2018, 2019). Using the merger rate in Eq. (5), Fig. 3 excludes light PBHs of masses $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$ as the totality of the DM. The mass range is further constrained, with the upper bound $m_{\rm PBH}\lesssim 10^{10}$ g from considerations over the effects of PBH evaporation at the onset of Big-Bang nucleosynthesis (BBN, blue contour) Carr et al. (2010). Also shown is the value $m_{\rm PBH}\sim 3\times 10^{3}$ g, corresponding to PBHs evaporating today when $k=1$ in Eq. (2) (vertical dashed line). For $k\lesssim 1$, constraints have been placed on PBH masses up to approximately $m_{\rm PBH}\sim 10^{5}$ g from the emission of photons Alexandre et al. (2024); Thoss et al. (2024) or neutrinos Chianese et al. (2024). However, the signal following the merger considered here is independent of $k$, offering a novel and complementary signature that does not depend on the specific model building details of the stabilized phase. Our results suggest that the window where memory-burdened PBHs constitute the DM is in tension with current measurements on the neutrino flux. However, this enhancement stems from limited theoretical understanding of non-Gaussian effects in PBH clustering and tidal disruption of PBH halos Delos et al. (2024). The bounds would be relaxed if the memory burden effect sets in earlier than the timescale $\tau_{\rm sc}$, resulting in a larger remnant mass Dvali et al. (2020, 2024a).
Figure 3: The DM fraction in PBHs of mass $m_{\rm PBH}$ as a function of the PBH mass. The results in this work consist in the bound from IceCube. Also shown are bounds from BBN considerations. In deriving the bound, we assumed that the memory burden effect occurs when the PBH has lost half of its initial mass. Different mass loss values are possible, potentially altering our findings as discussed in the text.
The results in Fig. 3 suggest that PBHs could make up a sizeable fraction of the DM in two mass ranges: $m_{\rm PBH}\lesssim 10^{5}$ g and $10^{8}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$. The lower mass range can be probed through gamma-ray flux measurements from experiments such as Auger (light green) Aab et al. (2020); Abreu et al. (2021a, b), Telescope Array (cyan) Abu-Zayyad et al. (2013); Ivanov (2020), and the Large High Altitude Air Shower Observatory (LHAASO) Addazi et al. (2022), which can detect gamma-rays above about 50 TeV. Studies of gamma and cosmic ray emissions from PBH clusters could further constraints $f_{\rm PBH}$, particularly for masses between $10^{3}\textrm{--}10^{5}$ g, while cosmic ray anisotropies might reveal nearby PBH subhalos Abbasi et al. (2010, 2011); Abeysekara et al. (2014, 2019); Amenomori et al. (2021), though more data is needed Giacinti and Sigl (2012). For higher-mass PBHs, upcoming neutrino observatories like IceCube-Gen2 radio array Abbasi et al. (2021b), the Radio Neutrino Observatory in Greenland (RNO-G) Aguilar et al. (2021), and the Giant Radio Array for Neutrino Detection (GRAND) Álvarez-Muñiz et al. (2020), will probe neutrinos at energies above 10 PeV with improved sensitivity. These observatories could constrain $f_{\rm PBH}$ by analyzing neutrino fluxes, especially when combined with data from Galactic sources, improving detection reach by $\mathcal{O}(10)$ over 10 years. Directional neutrino measurements will be crucial for distinguishing PBH contributions, providing a test for this scenario. Evaporating PBHs may emit gravitons, peaking at frequencies $\gtrsim 10^{30}$ Hz for $m_{\rm PBH}\lesssim 10^{8}$ g, beyond the reach of current GW detectors Domcke et al. (2022); Schmieden and Schott (2024); Gatti et al. (2024). GW abundance estimates suggest $\Omega_{\rm GW}\lesssim 10^{-12}$ at the peak.
Conclusions — In this Letter, we have explored the astrophysical implications to a PBH population affected by the memory burden. Given the halt in the evaporation process, light PBHs might live over cosmological timescales as the DM, undergoing binary merging and forming new BHs of similar mass at much smaller redshift than their formation epochs. We argued that newly-formed BHs initially do not carry memory burden and would then emit energetic particles in the form of cosmic rays, gamma rays, and neutrinos, which can be searched at dedicated facilities. We have estimated the integral spectrum in gamma rays from Galactic mergers and the differential spectrum of all-flavor neutrinos. Based on the current measurements by IceCube, we have derived the bounds on the PBH fraction, showing the tension with obtaining PBHs as the DM in the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$. If the merger rate is increased by primordial non-Gaussianity, the mass range that is potentially excluded sensibly widens. Future detectors will easily reach this signal. Other future work in this direction includes the careful study of the model with cosmic-rays from Galactic substructures.
Acknowledgements.
We thank Gia Dvali for useful discussions and for reading the manuscript. We further thank Gwenael Giacinti for useful comments. Finally, we are grateful to Concha Gonzalez-Garcia for correcting a previous mistake in the detectability of the neutrino mass nature at IceCube. We acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No. 12350610240 “Astrophysical Axion Laboratories”, as well as hospitality by the Istituto Nazionale di Fisica Nucleare (INFN).

### v4

@@
 Abstract
-The memory burden effect, stating that the amount of information stored within a system contributes to its stabilization, is particularly significant in systems with a high capacity for information storage such as black holes (BHs).
-In these systems, the evaporation process is halted, at the latest, after approximately half of the BH’s initial mass has been radiated away.
-Consequently, light primordial BHs (PBHs) of mass $m_{\rm PBH}\lesssim 10^{15}$ g, which are expected to have fully evaporated by present time, may remain viable candidates for dark matter (DM).
-In this scenario, we demonstrate that their mergers would continue to occur today, leading to the formation of “young” BHs that resume evaporating, producing ultrahigh-energy cosmic rays detectable by current experiments.
-The emission spectrum would be thermal in all Standard Model particle species, offering a clear and distinguishable signature.
-Current measurements of the isotropic neutrino flux at Earth are in tension with light PBHs as DM candidates within the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$.
-Contrarily to existing bounds, our results depend uniquely on the mass of the remnant and not on the specific model building details of the stabilized phase.
-We also discuss the potential for refining these constraints through gamma-ray and cosmic-ray observations, as well as gravitational wave detections.
-Introduction — Black holes (BHs), predicted by general relativity, are not only fascinating astrophysical objects but also intriguing sources of particle emission.
-A BH of mass $m_{\rm BH}$ emits high-energy particles lighter than its Hawking temperature, $T_{H}=(8\pi Gm_{\rm BH})^{-1}$ Hawking (1975).
+The memory burden effect, the idea that the amount of information stored within a system contributes to its stabilization, is particularly relevant for systems with a large information storage capacity, such as black holes.
+In these objects, the evaporation process halts, at the latest, once approximately half of the initial mass has been radiated away.
+As a result, light primordial black holes (PBHs) with mass $m_{\rm PBH}\lesssim 10^{15}\,\mathrm{g}$, which are traditionally assumed to have fully evaporated by the present time, may instead survive and constitute viable dark matter candidates.
+Ongoing mergers of such PBHs would give rise to “young” black holes that resume their evaporation, emitting ultrahigh-energy particles potentially detectable by current experiments.
+The resulting emission spectrum would be thermal across all Standard Model particle species, offering a clear and distinctive signature.
+We demonstrate that, if the memory burden effect activates after PBHs have lost around half of their initial mass, current measurements of the neutrino flux at Earth place strong constraints on such dark matter candidates for $m_{\rm PBH}\lesssim 10^{9}\,\mathrm{g}$.
+This suggests that the memory burden must set in at earlier stages of evaporation.
+Unlike existing bounds, our results depend solely on the mass of the remnant, and not on model-dependent details of the stabilized phase.
+We also discuss the potential for refining these constraints through observations of gamma rays, cosmic rays, and gravitational waves.
+I Introduction
+Black holes (BHs), predicted by general relativity, are not only fascinating astrophysical objects but also intriguing sources of particle emission.
+A BH of mass $m_{\rm BH}$ emits high-energy particles lighter than its Hawking temperature, $T_{\rm H}=(8\pi Gm_{\rm BH})^{-1}$ Hawking (1975).
 In the semiclassical regime where quantum gravity effects are neglected throughout their evaporation, the mass loss rate for an emitting BH follows $\dot{m}_{\rm BH}\propto m_{\rm BH}^{-2}$.
@@
 (2021); Auffinger (2023); Baker and Thamm (2023).
-However, it has been suggested that the information contained within a BH affects its decay, slowing down the evaporation process through a “memory burden” effect Dvali (2018a), which becomes significant after a portion of its mass has evaporated Dvali et al.
-(2020).
-The timescale that signals the breakdown of the semiclassical description, which is also the latest the memory burden effect sets in, reads
-$\tau_{\rm sc}\simeq(1-\epsilon^{3})\,\tau_{\rm evap}\,,$ (1)
-where $\tau_{\rm evap}$ is the BH evaporation time in the semiclassical limit and the prefactor accounts for a fraction $1-\epsilon$ of the BH mass being lost in the process.
-Here, we fix $\epsilon=1/2$.
+However, it has been suggested that the information contained within a BH affects its decay, slowing down the evaporation process through a “memory burden” effect Dvali (2018a); Dvali et al.
+(2020, 2024a), which becomes significant after a portion of its mass has evaporated Dvali et al.
+(2020, 2024a).
+The timescale that signals the breakdown of the semiclassical description and the set in of memory burden is
+$\tau_{\rm sc}\simeq q\,\tau_{\rm evap}\,,$ (1)
+where $\tau_{\rm evap}$ is the BH evaporation time according to the semiclassical rate and the prefactor accounts for the fraction $q\doteq\Delta m_{\rm PBH}/m_{\rm PBH}$ of the BH mass being lost during the semiclassical emission phase.
+Here, we fix $q=1/2$, corresponding to the latest time that memory burden may kick in.
 The subsequent dynamics of the memory-burdened BH for times $t\gtrsim\tau_{\rm sc}$ remain uncertain due to the influence of quantum gravity effects Dvali et al.
-(2020).
+(2020, 2024a).
 One possible outcome is that the configuration becomes unstable and eventually breaks down.
 Alternatively, evaporation could persist at a reduced rate, controlled by powers of the gravitational coupling, $\alpha_{\rm gr}\simeq S^{-1}$.
-In what follows, we assume the former scenario as the likely outcome indicated by theoretical analyses Dvali et al.
+In what follows, we assume the former scenario as the likely outcome which is further indicated by theoretical analyses Dvali et al.
 (2020, 2024a).
-When quantum gravity effects are taken into account, the BH evaporation rate is slowed down and the BH lifetime is extended as Dvali et al. (2020)
-$\tau\simeq\tau_{\rm sc}\,S^{1+k}\,,$ (2)
+When quantum effects are taken into account, the BH evaporation rate is slowed down and the BH lifetime is extended as Dvali et al. (2020)
+$\tau\simeq\tau_{\rm evap}\,S^{k}\,,$ (2)
 where $S=4\pi Gm_{\rm BH}^{2}$ represents the BH entropy.
-Theoretical models such as the holographic description of BHs Dvali (2018b) and the microscopic description in terms of marginally bound gravitons—the so-called BH $N$-portrait Dvali and Gomez (2013a, 2014a, 2014b)- independently suggest that $k=1$ Dvali et al.
-(2024a), while numerical analyses give $k\geq 1$ Dvali et al.
-(2024a).
+Theoretical models such as the holographic description of BHs Dvali (2018b) and the microscopic description in terms of marginally bound gravitons—the so-called BH $N$-portrait Dvali and Gomez (2013a, 2014a, 2014b)- independently suggest that $k=2$ Dvali et al.
+(2020, 2024a) (consistency requires $k\geq 1$).
 In what follows, we will derive bounds independent of this quantity.
@@
 (2020).
-For $k=1$, Eq.
-(2) predicts that memory-burdened BHs as light as $3\times 10^{3}\rm g$ can be DM candidates.
+For $k=2$, Eq.
+(2) implies that memory-burdened BHs as light as $\gtrsim 10^{3}\rm g$ can be DM candidates.
 A viable mechanism for their production within this mass range $3\times 10^{4}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$, here dubbed the ultralight BH window, was introduced in Ref.
@@
 (2021).
+[^1] See also Refs.
+Carr et al.
+(1994); Profumo (2025) for Planck-scale relics.
 A more precise mass range for memory-burdened PBHs constituting DM has been recently derived Alexandre et al.
@@
 (2025); Athron et al.
+(2025); Dvali et al.
 (2025).
-In this Letter, we discuss the observational consequences of memory-burdened PBHs as DM.
+In this paper, we discuss the observational consequences of memory-burdened PBHs as DM.
 We argue that these PBHs, locked in binary configurations, merge to form new BHs of similar mass.
@@
 Here, we work with natural units $\hbar=c=1$.
-Memory burden effect — Saturated configurations offer valuable insights into objects with high information-storage capacity, revealing connections between gravitational and non-gravitational systems.
+II Methods
+II.1 Memory burden effect
+Saturated configurations offer valuable insights into objects with high information-storage capacity, revealing connections between gravitational and non-gravitational systems.
 Notably, “saturons”, configurations of maximal entropy constrained by the unitarity bound Dvali (2021a), exemplify these objects.
 BHs are a prime example of saturated configurations, as they obey the entropy-area law Bekenstein (1973).
-Saturated configurations also arise in non-gravitational theories, such as renormalizable field theories Dvali (2021a, b, c).
-Saturated configurations not only follow an area law, but also share several properties with BHs, such as thermal-like evaporation Dvali and Sakhelashvili (2022); Dvali et al.
+Saturons also arise in non-gravitational theories, such as renormalizable field theories Dvali (2021a, b, c).
+These objects not only follow an area law, but also share several properties with BHs, such as thermal-like evaporation Dvali and Sakhelashvili (2022); Dvali et al.
 (2022a), the presence of an information horizon in the semiclassical regime Dvali (2021b, c, a, d); Dvali and Sakhelashvili (2022); Dvali et al.
@@
 (2022b, 2024b).
-Many features traditionally associated with BHs may arise from the broader behavior of saturated configurations, rather than being unique to gravity.
+As a consequence, many features traditionally associated with BHs may arise from the broader behavior of saturated configurations, rather than being unique to gravity.
 This framework provides new approaches for studying BHs from both theoretical and phenomenological perspectives.
 By analyzing the non-gravitational counterparts of BHs, we can use the calculability of renormalizable theories to better understand their enigmatic properties.
-Extending the study of saturated configurations to BHs allows us to predict new phenomenological signatures.
-For example, rotating saturons provide a microscopic description of maximally spinning BHs as graviton condensates with vorticity Dvali et al.
-(2022b), offering a topological explanation for the absence of Hawking radiation.
-This could lead to observable deviations in the gravitational radiation from BH mergers Dvali et al.
-(2024b).
-Furthermore, the memory burden effect, studied in calculable saturated configurations, is expected to manifest in BHs as well Dvali et al.
+Moreover, the study of this class of objects, allows for the extension of novel properties to BHs by universality.
+The memory burden phenomenon is a necessary feature of all saturons, including BHs.
+This stems from the requirement of generating a large microstate degeneracy to account for the entropy-area law.
+Achieving this requires a large number of modes with zero energy gap or “gapless”, localized on the support of the BH.
+Outside the object, these same degrees of freedom become gapped.
+As a result, for the BH to release its memory, an energy barrier needs to be overcome.
+Hawking radiation cannot do this due to its softness.
+The very same energy barrier is the one the BH eventually has to face as Hawking evaporation reduces its mass.
+We shall not rederive the memory burden mechanism here, and instead refer the reader to the relevant literature Dvali (2021a, b, c).
+Our focus is on features relevant for the merger process.
+The memory stored in a BH can be represented as $M$ qubits, which, due to their gaplessness, leads to a number of degenerate microstates $n_{\rm st}=2^{M}$, and therefore an entropy $S=\log n_{\rm st}\simeq M$.
+The amount of memory stored within the BH, corresponding to the total number of occupied memory modes $N_{\rm G}$, is determined statistically, according to the probability distribution Dvali et al.
+(2024a),
+[…]
+where $N_{\rm G}$ is effectively the order parameter responsible for the quantum backreaction Dvali et al.
+(2024a).
+On average, $N_{\rm G}\simeq S/2$.
+As BH decays, the increasing mass gap of the memory modes eventually halts the evaporation process.
+A careful dynamical analysis reveals that this transition is universally determined solely by the critical exponent, and can occur within a timescale $\tau_{\rm evap}/\sqrt{S}\lesssim\tau_{\rm SC}\lesssim\tau_{\rm evap}/2$ Dvali et al.
 (2020, 2024a).
-The memory stored in a BH can be represented as $M$ qubits, corresponding to the occupation number in the flavor space of Goldstone bosons arising from broken symmetries Dvali (2021a, b, c).
-The number of degenerate states is $n_{\rm st}=2^{M}$, leading to an entropy $S=\log n_{\rm st}\simeq M$.
-As BHs decay, the increasing mass gap of the Goldstone bosons halts evaporation by the half-mass decay time.
 [^2] The memory burden effect is an explicit realization of the so-called quantum-breaking of classical configurations Dvali et al.
 (2013), discussed within the framework of the BH $N$-portrait Dvali and Gomez (2013b, 2014b), which questions the validity of the classical description of BHs and finds that it is violated around $\tau_{\rm sc}$.
-This can be modeled using the probability distribution for memory-occupied modes Dvali et al.
-(2024a),
-[…]
-where $N_{\rm G}$ represents the order parameter setting the timescale of quantum backreaction Dvali et al.
-(2024a).
 The probability distribution in (3) leads to a natural spread in the mass distribution of PBHs Dvali et al.
 (2024a), offering a unique signature and further insights into BH evolution and mergers.
-A careful dynamical analysis reveals that an evaporating BH asymptotically evolves into a memory-burdened BH with around half of its initial mass Dvali et al.
-(2024a).
 While it is possible for a BH with empty memory to exist, allowing for its complete evaporation without memory burden, such cases are exponentially rare, as indicated by the probability distribution in Eq.
 (3).
-As suggested by the study of mergers Dvali et al.
+When two memory-burdened BHs merge, we expect a “young” BH to form, due to the large phase space volume occupied by these configurations.
+In fact, semiclassical BHs are realized when the memory modes are gapless and cost zero energy.
+At this point, the newly formed object possesses a large microstate degeneracy, and all such configurations need to be included in the amplitude, effectively saturating it Dvali (2021a).
+This saturation does not occur if one considers memory burdened BHs as the outcome, since BHs with different macroscopic quantum hairs have different energies and therefore do not saturate the phase space Dvali et al.
+(2024a).
+Equivalently stated, the formation of an initially classical BH is largely insensitive to the initial pre-merger configuration, due to the large entropy involved.
+A natural question is whether memory burden could dynamically induce significant deviations in the mass of the final object, compared to the initial configuration.
+The answer appears to be negative.
+In fact, as suggested by studies of mergers Dvali et al.
 (2024b) of memory-burdened configurations Dvali et al.
-(2024a), only the merging point is affected by the memory burden, potentially emitting additional radiation, while the merged object is unaffected.
-Mapped to the gravitational counterpart, these results imply that the memory burden of the two merging BHs leads to significant deviations in the gravitational wave signal only at ringdown, when a new classical BH forms with a memory pattern governed by Eq.
+(2024a), only the merging point is altered by the memory modes, and can lead to the radiation of a small mass fraction.
+This, however, may result in a final BH with a spin configuration that significantly differs compared to the case without memory.
+Mapped to the gravitational counterparts, these results suggest that the memory burden of the two merging BHs may source significant deviations in the gravitational wave signal during the ringdown phase, with the resulting BH showing a different angular momentum than classically expected.
+Since Hawking radiation after the merger is sensitive to the rotation of the object, which is also determined by the initial unknown spin of the two progenitors, we restrict our analysis to non-spinning BHs.
+The inclusion of rotational effects would strengthen our results.
+Although the merger dynamics can also lead to a partial release of the information, the newly formed BH has an area, and thus an entropy, approximately four times larger than the combined areas of the two initial BHs, providing considerable memory-storage capacity for all the initial information.
+Therefore, we expect that a new classical BH forms with a memory pattern governed by Eq.
 (3).
-Besides, the newly formed BH has an area, and thus an entropy, approximately four times larger than the combined areas of the two initial BHs, providing considerable memory to store all the non-emitted information.
-Therefore, post-merger the BH emits semiclassically for a time $\tau_{\rm sc}$ before stabilizing again.
+As a consequence, post-merger the BH emits semiclassically for a time $\tau_{\rm sc}$ before stabilizing again.
 [^3] The merger of extremal ultralight PBHs was discussed in Ref.
 Bai and Orlofsky (2020) where it was argued that Hawking emission is reactivated.
-Indeed, the memory burden effect shares similarities with BHs that are semiclassically stabilized by extremality Dvali et al.
+Indeed, memory-burdened PBHs may share some phenomenological similarities with BHs that are semiclassically stabilized by extremality Dvali et al.
 (2024a).
 This leads to important phenomenological consequences, which are discussed next.
-The cosmic ray flux from evaporating PBHs — The present fraction of memory-burdened PBHs formed at temperature $T_{f}$ is related to their primordial abundance $\beta$ through the conservation of the PBH comoving number, as
+II.2 High-energy particles from evaporating PBHs
+The present fraction of memory-burdened PBHs formed at temperature $T_{f}$ is related to their primordial abundance $\beta$ through the conservation of the PBH comoving number, as
 $f_{\rm PBH}\approx 10^{21}\,\beta^{\prime}\,\left(10^{8}{\rm\,g}/m_{\rm PBH}\right)^{1/2}\,.$ (4)
@@
 Here, we recast these bounds in terms of $f_{\rm PBH}$, reflecting the persistence of PBHs to present day.
-The differential rate of PBH binary mergers describes a distribution of PBHs that decoupled from the Hubble flow before matter-radiation equality Ali-Haïmoud et al.
+The differential rate of PBH binary mergers describes a distribution of PBHs that decoupled from the Hubble flow before matter-radiation equality Ali-Ha¨ımoud et al.
 (2017); Raidal et al.
@@
 (2021); Jedamzik (2020); Young and Hamers (2020); Jedamzik (2021)
-$R_{\rm PBH}(t)=\frac{4.9\times 10^{16}}{\rm kpc^{3}\,yr}\,f_{\rm PBH}^{\frac{53}{37}}\,\left(\frac{t_{0}}{t}\right)^{\frac{34}{37}}\,\left(\frac{m_{\rm PBH}}{\rm g}\right)^{-\frac{32}{37}}S\,,$ (5)
+$R_{\rm PBH}(t)\simeq\frac{5.7\times 10^{-66}}{\rm cm^{3}\,s}f_{\rm PBH}^{\frac{53}{37}}\left(\frac{t_{0}}{t}\right)^{\frac{34}{37}}\left(\frac{2m_{\rm PBH}}{10^{10}\rm g}\right)^{-\frac{32}{37}}S\,,$ (5)
 where the suppression factor $S$ is the product of two components: a term $S_{1}\approx 0.24$ which accounts for the interactions between the binary system and the surrounding DM inhomogeneities, as well as neighboring BHs Hütsi et al.
@@
 Here, $x\equiv\left(t/t_{0}\right)^{0.44}f_{\mathrm{PBH}}$.
-The function $S_{2}(x)$ attains a minimum at $10^{-2}$, consistently with numerical simulations Inman and Ali-Haïmoud (2019); Franciolini et al.
+The function $S_{2}(x)$ attains a minimum at $10^{-2}$, consistently with numerical simulations Inman and Ali-Ha¨ımoud (2019); Franciolini et al.
 (2022).
@@
 […]
+[…]
 The emission of newly formed BHs in Eq.
@@
 Averaging the differential flux over the angular direction $\theta$ and integrating in the energy bins yields the integral flux $\Phi_{i}^{\rm gal}(E)$.
-The extra-galactic (eg) contribution to the differential flux of the species $i$ is
+The extra-galactic (eg) contribution to the differential flux of the species $i$ is [^5] We thank Paolo Panci for noticing a typo in the original formula.
 $\frac{{\rm d}\Phi_{i}}{{\rm d}E}\bigg{|}_{\rm eg}\simeq\frac{\tau_{\rm sc}}{4\pi}\int_{0}^{z_{f}}\!{\rm d}z\left|\frac{{\rm d}t}{{\rm d}z}\right|R_{\rm PBH}(t(z))\frac{{\rm d^{2}}N_{i}(E^{\prime})}{{\rm d}E\,{\rm d}t}e^{-\eta_{i}(z)}\,,$ (7)
@@
 The opacity $\eta_{i}$ describes the attenuation between $z$ and today at energy $E$.
-Results — Figure 1 shows the integrated gamma-ray flux $\Phi_{i}^{\rm gal}(E)$ in Eq.
+More details can be found in Dondarini et al.
+(2025).
+III Results
+Figure 1 shows the integrated gamma-ray flux $\Phi_{i}^{\rm gal}(E)$ in Eq.
 (6), produced by newly-formed BHs from PBH mergers and for $f_{\rm PBH}=1$, for different values of the PBH masses and the conservative merger rate in Eq.
 (5).
-We have not included the eg contribution since, at energies above the TeV scale, the Universe is opaque to gamma rays due to gamma-gamma absorption (see Cao et al.
-(2023) for a review).
+The absorption of gamma rays over cosmological timescales, as well as cascade effects, are implemented following Ref.
+Dondarini et al.
+(2025).
 Results are compared with the upper limits at 95% confidence level (CL) from the Pierre Auger Observatory (Auger) Savina (2021); Abreu et al.
@@
 Figure 1: The integrated gamma-ray flux emitted by newly-formed BHs from memory-burdened Galactic PBH mergers, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$.
+$m_{i}$ denotes the mass of the PBHs post-merger which is roughly twice the mass of the DM.
 Also shown are the upper limits at 95% CL from Auger Savina (2021); Abreu et al.
@@
 (2017).
-We now discuss the emission of energetic neutrinos of all flavors by the same sources.
+We now discuss the emission of energetic neutrinos by the same sources.
 The differential neutrino flux per solid angle and energy is reported by the IceCube collaboration using the cascades collected over 7.5 years of operation at energies above 60 TeV Abbasi et al.
@@
 (2020).
-Above the PeV, an upper limit at 95% CL on the neutrino flux is reported Aartsen et al.
-(2018, 2019).
+Below around 10 PeV, the neutrino flux is bound by observations from ANTARES Albert et al.
+(2024).
+At energies above 10 PeV, an upper limit at 95% CL on the neutrino flux is reported by the IceCube High-Energy Starting Events (HESE) Aartsen et al.
+(2018, 2019) and by Auger Abdul Halim et al.
+(2023).
+IceCube reports candidate detections through the Northern Sky Tracks (NST) Abbasi et al.
+(2022) and HESE Abbasi et al.
+(2021a).
 In Fig.
-2 we compare these experimental results with the the differential flux of all-flavor neutrinos sourced by newly-formed BHs.
-We include the Galactic (continuous) and extragalactic (dashed) contributions assuming the merger rate in Eq.
+2 we show these measurements, together with the recent high-energy event detected by KM3NeT collaboration Aiello et al.
+(2025) as well as the Glashow resonant observed by IceCube Aartsen et al.
 (2021).
-Figure 2: The differential flux of all-flavor neutrinos sourced by newly-formed BHs, for a monochromatic PBHs distribution with $f_{\rm PBH}=1$.
-We show the Galactic (continuous) and the extra-galactic (dashed) contributions.
-The theoretical flux is compared with the all-flavor neutrino flux above 60 TeV Abbasi et al.
-(2021a); Aartsen et al.
-(2020) and with the upper bound for ultra-high energy neutrinos Aartsen et al.
-(2018, 2019).
-Discussion — Assuming that a fraction $f_{\rm PBH}$ of the DM relic abundance consists of light PBHs, we derive a constraint on the model in the $(m_{\rm PBH},f_{\rm PBH})$ plane by performing a least-square analysis that accounts for the neutrino flux measurements and upper bounds reported by the IceCube collaboration Abbasi et al.
+We compare these experimental results with the per-flavor neutrinos differential flux sourced by merged PBHs.
+The Galactic (dashed), extragalactic (dot-dashed), and total (solid) contributions are obtained assuming the merger rate in Eq.
+(5).
+Figure 2: Per-flavor neutrino fluxes for various monochromatic PBH mass distributions with $f_{\rm PBH}=1$, normalized by a factor of 2 22 for better readability.
+Dashed (dot-dashed) lines represent the Galactic (extra-galactic) contributions, while solid lines show their sum.
+The black line in the upper right corner indicates the current upper limit from IceCube High-Energy Starting Events (HESE) Aartsen et al.
+(2018), while the one on the left shows limits from ANTARES Albert et al.
+(2024).
+The gray line corresponds to the upper bound set by Auger Abdul Halim et al.
+(2023).
+Black (blue) points indicate IceCube detections from Northern Sky Tracks (NST) Abbasi et al.
+(2022) and HESE Abbasi et al.
+(2021a), respectively.
+The green cross marks the Glashow resonance event observed by IceCube Aartsen et al.
+(2021), and the red point shows the high-energy neutrino KM3-230213A detected by KM3NeT Aiello et al.
+(2025).
+IV Discussion
+Assuming that a fraction $f_{\rm PBH}$ of the DM relic abundance consists of light PBHs, we derive a constraint on the model in the $(m_{\rm PBH},f_{\rm PBH})$ plane by performing a least-square analysis that accounts for the neutrino flux measurements and upper bounds reported by the IceCube collaboration Abbasi et al.
 (2021a); Aartsen et al.
@@
 (5), Fig.
-3 excludes light PBHs of masses $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$ as the totality of the DM.
+3 excludes light PBHs of masses $m_{\rm PBH}\lesssim 10^{9}\,\rm g$ as the totality of the DM (filled cyan area).
 The mass range is further constrained, with the upper bound $m_{\rm PBH}\lesssim 10^{10}$ g from considerations over the effects of PBH evaporation at the onset of Big-Bang nucleosynthesis (BBN, blue contour) Carr et al.
+(2010, 2021c); Thoss and Burkert (2025).
+Also shown is the value $m_{\rm PBH}\sim 3\times 10^{3}$ g (vertical dashed line), for which $\tau_{\rm sc}\simeq t_{0}$ when $k=2$, c.f., Eq.
 (2).
-Also shown is the value $m_{\rm PBH}\sim 3\times 10^{3}$ g, corresponding to PBHs evaporating today when $k=1$ in Eq.
-(2) (vertical dashed line).
-For $k\lesssim 1$, constraints have been placed on PBH masses up to approximately $m_{\rm PBH}\sim 10^{5}$ g from the emission of photons Alexandre et al.
+For $k=2$, constraints have been placed on PBH masses up to approximately $m_{\rm PBH}\sim 10^{5}$ g from the emission of photons Alexandre et al.
 (2024); Thoss et al.
 (2024) or neutrinos Chianese et al.
-(2024).
-However, the signal following the merger considered here is independent of $k$, offering a novel and complementary signature that does not depend on the specific model building details of the stabilized phase.
+(2025) (see Dondarini et al.
+(2025) for a complete analysis).
+However, the signal following the merger considered here is not sensitive to $k$, offering a novel and complementary signature that does not depend on the specific details of the stabilized phase.
 Our results suggest that the window where memory-burdened PBHs constitute the DM is in tension with current measurements on the neutrino flux.
-However, this enhancement stems from limited theoretical understanding of non-Gaussian effects in PBH clustering and tidal disruption of PBH halos Delos et al.
-(2024).
-The bounds would be relaxed if the memory burden effect sets in earlier than the timescale $\tau_{\rm sc}$, resulting in a larger remnant mass Dvali et al.
-(2020, 2024a).
 Figure 3: The DM fraction in PBHs of mass $m_{\rm PBH}$ as a function of the PBH mass.
-The results in this work consist in the bound from IceCube.
-Also shown are bounds from BBN considerations.
-In deriving the bound, we assumed that the memory burden effect occurs when the PBH has lost half of its initial mass.
-Different mass loss values are possible, potentially altering our findings as discussed in the text.
-The results in Fig.
-3 suggest that PBHs could make up a sizeable fraction of the DM in two mass ranges: $m_{\rm PBH}\lesssim 10^{5}$ g and $10^{8}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$.
-The lower mass range can be probed through gamma-ray flux measurements from experiments such as Auger (light green) Aab et al.
-(2020); Abreu et al.
-(2021a, b), Telescope Array (cyan) Abu-Zayyad et al.
-(2013); Ivanov (2020), and the Large High Altitude Air Shower Observatory (LHAASO) Addazi et al.
-(2022), which can detect gamma-rays above about 50 TeV.
-Studies of gamma and cosmic ray emissions from PBH clusters could further constraints $f_{\rm PBH}$, particularly for masses between $10^{3}\textrm{--}10^{5}$ g, while cosmic ray anisotropies might reveal nearby PBH subhalos Abbasi et al.
+The results in this work consist in the bound from IceCube (filled cyan area).
+Also shown are bounds from BBN considerations, taken from Carr et al.
+(2010, 2021c); Thoss and Burkert (2025).
+In deriving the bound, we assume that the memory burden effect occurs when the PBH has lost half of its initial mass.
+Different mass-loss fraction are possible, potentially altering our findings as discussed in the text.
+As shown in Fig.
+3, only PBHs in the mass range $10^{9}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{10}$ could make up a sizable fraction of the DM.
+Obviously, it is possible that memory burden sets in before half of the mass has been emitted semi-classically.
+This possibility is actually supported in theoretical studies, where it has been found that memory burden may emerge after an $\mathcal{O}(1/\sqrt{S})$ fraction of the initial mass has been emitted Dvali (2018a); Dvali et al.
+(2019, 2020); Alexandre et al.
+(2024); Dvali et al.
+(2024a).
+Independently motivated studies found that, on comparable timescales, the semiclassical description of BHs ceases to be accurate Dvali and Gomez (2013a, b, 2014a, 2012); Dvali et al.
+(2013); Dvali and Panchenko (2015); Michel and Zell (2023).
+In such scenario, the bounds presented here would no longer apply.
+Comparable, and potentially stronger constraints are expected in a similar mass range due to gamma-ray flux measurements by the Large High Altitude Air Shower Observatory (LHAASO) Addazi et al.
+(2022), which detected gamma-rays above about 10 TeV.
+Studies of gamma and cosmic ray emissions from PBH clusters could further constraints $f_{\rm PBH}$, particularly for masses smaller than $10^{5}$ g, while cosmic ray anisotropies might reveal nearby PBH subhalos Abbasi et al.
 (2010, 2011); Abeysekara et al.
@@
 (2020), will probe neutrinos at energies above 10 PeV with improved sensitivity.
+In this direction, we ought to mention the recent measurement of the high-energy - above $10^{2}\,$eV - neutrino event detected by the KM3NeT collaboration Aiello et al.
+(2025).
 These observatories could constrain $f_{\rm PBH}$ by analyzing neutrino fluxes, especially when combined with data from Galactic sources, improving detection reach by $\mathcal{O}(10)$ over 10 years.
@@
 GW abundance estimates suggest $\Omega_{\rm GW}\lesssim 10^{-12}$ at the peak.
-Conclusions — In this Letter, we have explored the astrophysical implications to a PBH population affected by the memory burden.
-Given the halt in the evaporation process, light PBHs might live over cosmological timescales as the DM, undergoing binary merging and forming new BHs of similar mass at much smaller redshift than their formation epochs.
-We argued that newly-formed BHs initially do not carry memory burden and would then emit energetic particles in the form of cosmic rays, gamma rays, and neutrinos, which can be searched at dedicated facilities.
-We have estimated the integral spectrum in gamma rays from Galactic mergers and the differential spectrum of all-flavor neutrinos.
-Based on the current measurements by IceCube, we have derived the bounds on the PBH fraction, showing the tension with obtaining PBHs as the DM in the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$.
-If the merger rate is increased by primordial non-Gaussianity, the mass range that is potentially excluded sensibly widens.
-Future detectors will easily reach this signal.
-Other future work in this direction includes the careful study of the model with cosmic-rays from Galactic substructures.
+V Conclusions
+In this paper, we have explored the astrophysical implications to a PBH population affected by the memory burden effect.
+Due to the associated halt in the evaporation process, light PBHs might live over cosmological timescales and constitute viable DM candidates, undergoing binary merging and forming new BHs of similar mass at much later cosmic times than their formation epochs.
+We argued that newly-formed BHs initially do not carry memory burden and can therefore emit energetic particles, in the form of cosmic rays, gamma rays, and neutrinos, which can be searched at dedicated facilities.
+We have estimated the integral spectrum in gamma rays from Galactic mergers and the differential spectrum of per-flavor neutrinos.
+Based on the current available measurements by IceCube, we have derived bounds on the PBH fraction.
+We found that, for a late memory burden onset, after the BH has lost roughly half of its initial mass, PBHs can account for the entirety of DM only in a small window $10^{9}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{10}$.
+Our findings are weakened if the memory burden sets in at parametrically earlier times.
+However, if the merger rate is increased by primordial non-Gaussianity, the mass range that is potentially excluded sensibly widens and our constraints strengthen.
+Future detectors will be able to probe even deeper into the remaining viable parameter space.
+Additional future directions include the careful study of the model with cosmic-ray and gamma-ray signatures from Galactic substructures.
+Note Added.
+After the completion of this work, Ref.
+Dondarini et al.
+(2025) performed an in depth analysis of neutrinos, gamma-rays and early Universe signatures (CMB) resulting from the presented mechanism.
+We refer the interested reader to  Dondarini et al.
+(2025) for a detailed discussion of the existing constraints as well as the characterization of the duration of the semiclassical emission phase (given by $q$ in (1)) that allows $f_{\rm PBH}=1$ in the entire mass window.
+Our result in Fig.
+3 is qualitatively consistent to theirs.
+Some minor discrepancies follow from the fact that Dondarini et al.
+(2025) adopted slightly different datasets and included the energy binning of the experimental points in the analysis.
 Acknowledgements.
@@
 We further thank Gwenael Giacinti for useful comments.
-Finally, we are grateful to Concha Gonzalez-Garcia for correcting a previous mistake in the detectability of the neutrino mass nature at IceCube.
+Finally, we are grateful to Concha Gonzalez-Garcia for correcting a previous mistake in the detectability of the neutrino mass nature at IceCube and for further discussions.
+MZ is thankful to Alessandro Dondarini, Giulio Marino and Paolo Panci for useful ongoing discussion on related topics.
 We acknowledge support by the National Natural Science Foundation of China (NSFC) through the grant No.
-12350610240 “Astrophysical Axion Laboratories”, as well as hospitality by the Istituto Nazionale di Fisica Nucleare (INFN).
+12350610240 “Astrophysical Axion Laboratories”.
+LV acknowledges support by the Istituto Nazionale di Fisica Nucleare (INFN) through the “QGSKY” Iniziativa Specifica project.
+LV thanks the Tsung-Dao Lee Institute for its hospitality during the final stages of this work.